# Anticipating Object State Changes

### v2

@@ Abstract
In this work, we introduce (a) the new problem of anticipating object state changes in images and videos during procedural activities, (b) new curated annotation data for object state change classification based on the Ego4D dataset, and (c) the first method for addressing this challenging problem. Solutions to this new task have important implications in vision-based scene understanding, automated monitoring systems, and action planning. The proposed novel framework predicts object state changes that will occur in the near future due to yet unseen human actions by integrating learned visual features that represent recent visual information with natural language (NLP) features that represent past object state changes and actions. Leveraging the extensive and challenging Ego4D dataset which provides a large-scale collection of first-person perspective videos across numerous interaction scenarios, we introduce an extension noted Ego4D-OSCA that provides new curated annotation data for the object state change anticipation task (OSCA). An extensive experimental evaluation is presented demonstrating the proposed method’s efficacy in predicting object state changes in dynamic scenarios. The performance of the proposed approach also underscores the potential of integrating video and linguistic cues to enhance the predictive performance of video understanding systems and lays the groundwork for future research on the new task of object state change anticipation. The source code and the new annotation data (Ego4D-OSCA) will be made publicly available [^1] https://projects.ics.forth.gr/cvrl/osca/.
Figure 1: We introduce the new problem of anticipating object state changes in videos of procedural activities, noted OSCA. The decision point in the timeline is placed at the onset of the next anticipated, yet unobserved, human action. The objective is to accurately predict, at this point, the object state change class that will occur, if any, during the subsequent, yet unobserved action. This involves understanding the dynamics of past and current interactions and how they will affect the object’s state. An object state change (e.g. deform) refers to a functional and possibly to a physical change in an object’s attributes/properties. It is realized based on the transition from a pre-state (initial) to a post-state (final) occurring at the Point of No Return (PNR) time during an object state-modifying action (e.g. cut fish fillet using a knife).
@@ 1 Introduction
The problem of Object State Classification (OSC) is defined as the multi-class recognition of an object’s state in a still image [22, 17], or the initial and final object states in videos that demonstrate state-modifying actions [18]. The binary object state change classification variant is also related to the detection of state change occurrences in a short video clip [7, 18]. Researchers have only recently started to focus on methods for the representation and understanding of object state changes in videos in the context of state-modifying actions [38, 44, 41], which can also be seen as transformations [50]. Existing benchmarks largely ignore object state changes and focus on traditional types of annotations related to object type, location, or shape, attributes, affordances, and human actions.
@@
In this work, we take one step beyond the Object State Classification and the Action Anticipation tasks by introducing the new task of Object State Change Anticipation (OSCA) in videos. OSCA focuses on the multi-class prediction of the state change that will occur on an object during the yet unseen, next action in an instructional video. Specifically, as shown in Fig. 1, at a certain decision point in time that is at the start of the next, yet unobserved action, we aim to predict the object state change class that will occur. The object state change will occur at the “Point of No Return” (PNR) timestamp during the next action [18]. To tackle this new task, we introduce a new formulation leveraging on the cumulative history of the textual description of recognized preceding actions and object states, up to the decision point where this prediction is performed. The integration of the former information with visual information concerning the recent past is a key idea to model this historical context effectively. Our approach may also predict that no object state change will occur in the forthcoming action which implies the anticipation of a no-state-modifying action. We evaluate the proposed approach toward the newly proposed OSCA problem to establish baseline results for automated change state anticipation in long instructional videos and to also investigate the potential impact and effectiveness of the extracted information in other related vision-based anticipation tasks. We assess the performance of our proposed method for the OSCA problem and show initial results based on a proposed extension of the popular Ego4D video dataset. We hence build on the large-scale and challenging Ego4D dataset [18] which provides ego-centric videos by augmenting the available annotation data with labels for the object state changes based on the initial and the final object states for any state-modifying actions in a subset of videos related to the Hand & Object Interactions benchmark  [^2] https://ego4d-data.org/docs/benchmarks/hands-and-objects/. This results in the Ego4D-OSCA, a variant of the Ego4D dataset that will become available to the computer vision community.
Thus, the main contributions of this work are:
@@
Figure 2: Examples of modifying actions from the “deform” and “remove” object state change classes. Each example is represented by a triplet of frames (pre-state, PNR, post-state). A state change is associated with various actions taking place in diverse environments and scenarios, emphasizing the complexity and challenges introduced in the OSCA problem.
@@ 2 Related Work
Object states capture dynamic aspects of object appearance and/or functionality and are subject to visually perceivable changes, as a result/effect of state-modifying actions. They are also known as object fluents related to changeable object attributes [24, 2, 25]. Since there is no prior work on the problem of Object State Change Anticipation (OSCA) that we introduce, we review the literature on the closely related topics of action anticipation, next-active object anticipation, and object state classification in images as well as the interplay between object state estimation and action recognition in videos.
@@ 2.1 Object State Classification/Recognition
Object State Change Estimation & Action Recognition in Videos: Object state changes have been considered a meaningful information source in video-based human action understanding and recognition (HAR). In HAR, object state changes are often considered complementary attributes to the visual representation of actions. These changes are typically derived within the visual domain through the utilization of explicit models for object detection and state estimation [11, 41, 26], or indirect modeling of object states based on general scene changes resulting from action execution [1, 5, 4]. Several methods exploit object states implicitly to estimate the type of action performed. The work in [1] was among the first to propose a method to automatically discover object states and the associated manipulation actions from videos by leveraging a discriminative clustering framework that jointly models the temporal order of object states and manipulation actions. The work in [25] explored the recognition of object fluents (changeable attributes of objects) and tasks (goal-oriented human activities) in egocentric videos using a hierarchical model that represents tasks as concurrent and sequential object fluents. Moreover, [27] focuses on understanding human actions within videos by analyzing complex interactions across multiple interrelated objects by recognizing different state changes of these objects. In [43, 41] a multi-task self-supervised framework is proposed that allows the temporal localization of object state changes and state-modifying actions in uncurated web videos.
@@ 2.2 Action & Next-Active Object Anticipation
Figure 3: Two of the nine object state change classes that have been super-annotated in the proposed Ego4D-OSCA dataset, ”deposit” and ”remove,” are illustrated. Pre-/post-state labels for two state-modifying actions are shown as distinct video segments. ”Deposit” and ”remove” are inverse state changes, where the pre-deposit matches the post-remove, and the pre-remove matches the post-deposit, as indicated by frames and shapes of the same color.
@@ 3 Ego4D-OSCA Dataset
Dataset annotation: To enrich the LTA benchmark, we have integrated object state annotations extracted from the dataset as follows. The original Ego4D dataset does not include annotations for the specific state labels of individual video frames. Instead, annotations about state changes are provided, which relate to entire video segments. Additionally, the dataset includes annotations for bounding boxes and object classes across seven critical frames within each video segment. These frames are temporally centered around the occurrence of the state change that occurs within each video segment. Based on this information, we super-annotate certain critical frames of each video segment with state-related labels as follows. For each video segment, we annotate the initial and final frames as $pre\_X$ and $post\_X$, respectively, where $X$ denotes the label of the state change. Furthermore, in line with the semantic implications of these changes, we establish three pairs of state changes. Each pair is constructed under the premise that the first action is the inverse of the second concerning the resulting state change. For instance, if $X$ and $Y$ represent inverse state changes, then the labels $pre\_X$ and $post\_Y$ are considered samples of identical states. A similar correspondence applies between $pre\_Y$ and $post\_X$. For example, the states $pre\_remove$ and $post\_deposit$ are considered identical, since $remove$ and $deposit$ constitute a pair of inverse state changes. Figure 3 delineates the specifics for these two super-annotated state change classes. The same condition is true for the state change classes of activate-deactivate and construct-deconstruct. The full set of pre- and post- object state pairs that constitute the target set of object state changes appear in the supplementary material. It is important to note that there exists a subset of video segments containing actions that do not induce state changes and, therefore, these segments are not considered for annotation.
@@
Details on the annotation process: The annotation process for the state transitions is applied to the $pre\_Y$ and $post\_X$ frames in each video segment. Overall, the annotation process consists of the next 4 steps (a schematic representation of the annotation pipeline is shown in Figure 4). First, the PNR moment of the video segment being examined for annotation is compared to the PNR moment of the segment that has been previously annotated. If the PNR of the previously annotated segment is located after the PNR of the segment under examination, then the segment under examination is rejected. The reasoning behind this decision has to do with the learning of the segment features related to state transitions. This alignment of the two PNRs signifies that there is an overlap between the state transition actions of the two segments and therefore the feature learning becomes more challenging. Subsequently, it is examined if the object undergoing state change is occluded. If this is the case, the frame is rejected. Then, the bounding box area of the object is evaluated, and if it is below 100 square pixels—a threshold empirically chosen and commonly used in annotation tools like Voxel-51—the frame is discarded. Finally, the frame that has passed all the previous checks is annotated with the appropriate state label that pertains to the state transition action.
Dataset Statistics: The proposed Ego4D-OSCA dataset is compiled using a subset of the popular, large-scale Ego4D v2 dataset [18] that contains egocentric videos for a large variety of human daily living or work activities. In Table 1 we compare the proposed Ego4D-OSCA dataset with existing image and video datasets that also provide annotations related to object states. The target tasks performed using each of the datasets are also noted. Moreover, Figure 5 provides statistics about the state change classes. More dataset statistics can be found in the supplementary material.
Figure 5: Statistics for the Ego4D-OSCA dataset by object state change class, with 61.858 training and 31.846 testing clips.
@@ 4 Object State Change Anticipation - Baseline
The proposed framework, depicted in Fig. 6, draws inspiration from the efficacy of combining visual and lexical information for semantic action/activity encoding. To achieve this, it adopts a three-stream architecture. Within this design, a visual encoding module is tasked with capturing the visual attributes of ongoing actions, while two lexical-based encoders are employed to extract the semantic nuances from a procedural-oriented representation of past actions and object states. The framework fuses these distinct representations towards the unified objective of anticipating the next object state. This task entails the estimation of the forthcoming state in which the object of interest will reside during the subsequent action. By integrating visual and lexical cues, the framework tries to holistically capture the underlying dynamics and contextual intricacies governing object-state transformations across sequential actions.
@@
Action & State History Encoders: As illustrated in Fig. 6, both encoders exploit the model design of the lexical encoder of the VLMAH model [28], which follows a NLP neural network design consisting of BiLSTM and MLP components. The decision to employ a simple NN for encoding the history, instead of utilizing LLMs was motivated by several factors. Firstly, the computational efficiency of LLMs such as GPT-3 or LLaMA, often entails significant resource requirements for training and inference [48], whereas a simpler neural network architecture mitigates computational overhead. Secondly, LLMs are pre-trained on general text corpora and may not capture the domain-specific nuances inherent in the textual data related to action histories and object states. Additionally, the simplicity of the chosen architecture facilitates interpretability, data efficiency, and customization, affording greater control over the model’s behavior and adaptation to the task’s requirements.
@@
During training, the proposed framework leverages oracle action and state detectors to provide the action and state history, respectively, for each clip in the dataset (see Fig. 6). These detectors estimate the current action and object state observed in the clip, serving as ground truth annotations for training purposes. However, it is important to note that for inference toward real-world applications, there is a requirement for current action and object state recognition models to provide input to the framework. Consequently, our proposed model is solely tasked with the learning objective of next-object anticipation during training, focusing exclusively on predicting the future state of the object. By decoupling the training and inference phases in this manner, the model can effectively learn the dynamics of object-state transitions without the added complexity of simultaneously predicting the current action.
@@ 5 Implementation, Experiments and Results
Evaluation metrics: The evaluation of all examined models was conducted using top-1 and top-5 mean accuracy, following standard practices in the relevant literature.
@@ 5.1 OSCA Results
OSCA under actual action and state recognition: We conducted additional experiments where the oracle-based modules, namely the action recognizer and the current object-agnostic state estimator, were substituted with baseline models. For this experimental setting, the output of these models is utilized to populate the action and state histories, therefore aiding our OSCA model in estimating the anticipated state. We employed the following models: (a) the well-established SlowFast [12] model for action recognition; (b) the object-agnostic state classification method proposed by [17] as the object-agnostic current state classifier, with minor modifications, as described below.
@@
Action recognizer: Regarding the action classification module, we fine-tuned the SlowFast model [12] on a subset of the original Ego4D dataset. Specifically, we obtained the train/validation/test splits using the training set provided for the (LTA) long-term anticipation task of Ego4D based on the $60/20/20$ split scheme. The adaptation of the LTA data into the action recognition task resulted in $5754$ action classes and a total of $\approx 65K$ video clips (with a mean of $\approx 10.7$ sample clips per action class). SlowFast achieved $12.86\%$ Top-1 and $33.69\%$ Top-5 accuracy for the task of current action recognition. This low performance can be attributed to the extensive number of action classes and limited samples per class, as well as in Ego4D’s inherent motion and appearance similarities across different actions, e.g. take cup - take bottle, tie string - tie rope.
Model Top@1/5 mAcc
VID-A 23.93 / 89.10%
VNLP (O-Action) 25.59 / 83.06%
VNLP (Action [12]) 23.04 / 81.31%
VNLP (O-State) 34.35 / 85.92%
VNLP (State [17]) 30.23 / 91.80%
VNLP (O-Action, O-State) 35.60 / 88.14%
VNLP (Action [12] , State [17]) 28.97 / 92.25%
@@
Table 2: Object state change anticipation performance for various model configurations (O-: Oracle recognizer, VID-A: vision-only state anticipation model).
@@
Object state recognizer: Additionally, for the object-agnostic state history, we adapt the model of [17]. This model relies on the outputs of two distinct state classifiers. Each classifier receives the first (pre) or the last (post) frame of each video segment as input to predict the object state label for the respective frame. The prediction of the state change label for the video segment considers both outputs and is derived based on the following rules. If the object state predictions are $pre\_X$ and $post\_X$, respectively, the inferred state change for the video segment is denoted as $X$. Conversely, if the classifiers predict $pre\_X$ and $post\_Y$, where $X$ and $Y$ are distinct and represent inverse state changes, it is concluded that no state change has occurred. Finally, if neither of the conditions mentioned above is met, the state change prediction defaults to the output of the second classifier; i.e., if the prediction is $post\_Y$, the state change for the video segment is identified as Y. For example, if the predictions of the two classifiers are $pre\_activate$ and $post\_activate$ the prediction for the state change would be $activate$. Likewise, if the predictions are $pre\_activate$ and $post\_deactivate$ the prediction of the state change would be that of no change. This object-agnostic state recognizer showcased a mean frame-wise state recognition accuracy of $25.4\%$.
In our experiments when we replaced the oracles with the aforementioned realistic recognition models to populate the action and state history buffers that are being considered by the baseline OSCA model, we observed a significant drop in accuracy (rows 4, 6, & 8 of Table 2). This difference in accuracy underscores the critical role of precise current action and object state recognition in effectively anticipating future object states within dynamic environments.

### v3

@@ Abstract
In this work, we introduce (a) the new problem of anticipating object state changes in images and videos during procedural activities, (b) new curated annotation data for object state change classification based on the Ego4D dataset, and (c) the first method for addressing this challenging problem. Solutions to this new task have important implications in vision-based scene understanding, automated monitoring systems, and action planning. The proposed novel framework predicts object state changes that will occur in the near future due to yet unseen human actions by integrating learned visual features that represent recent visual information with natural language (NLP) features that represent past object state changes and actions. Leveraging the extensive and challenging Ego4D dataset which provides a large-scale collection of first-person perspective videos across numerous interaction scenarios, we introduce an extension noted Ego4D-OSCA that provides new curated annotation data for the object state change anticipation task (OSCA). An extensive experimental evaluation is presented demonstrating the proposed method’s efficacy in predicting object state changes in dynamic scenarios. The performance of the proposed approach also underscores the potential of integrating video and language cues to enhance the predictive performance of video understanding systems and lays the groundwork for future research on the new task of object state change anticipation. The source code and the new annotated data will be made publicly available [^1] https://projects.ics.forth.gr/cvrl/osca/.
Figure 1: We introduce the new problem of anticipating object state changes in videos of procedural activities, noted OSCA. The decision point in the timeline is placed at the onset of the next anticipated, yet unobserved action. The objective is to predict accurately, at this point, the object state change class that will occur, if any, during the subsequent, yet unobserved action. This involves understanding the dynamics of past and current interactions and how they will affect the object’s state. An object state change (e.g. deform) refers to a physical and possibly functional change in an object’s attributes/properties. It is realized based on the transition from a pre-state (initial) to a post-state (final) occurring at the Point of No Return (PNR) time during an object state-modifying action (e.g. cut fish fillet using a knife).
@@ 1 Introduction
Figure 2: Examples of modifying actions from the “deform” and “remove” object state change classes represented by a triplet of frames (pre-state, PNR, post-state). Each state change is associated with various actions occurring in diverse environments/scenarios, emphasizing the complexity and challenges introduced in the OSCA problem.
@@
The problem of Object State Classification (OSC) is defined as the multi-class recognition of an object’s state in a still image [23, 16], or the initial and the final object states in a video that demonstrates one or more state-modifying actions [19]. The binary object state change classification variant is also related to detecting state change occurrences in a short video clip [7, 19]. Researchers have only recently started to focus on methods for the representation and understanding of object state changes in videos in the context of state-modifying actions [41, 47, 44], which can also be seen as transformations [54]. Existing benchmarks largely ignore object state changes and focus on traditional types of annotations related to object type, location, or shape, attributes, affordances, and human actions.
In this work, we take one step beyond the Object State Classification and the Action Anticipation tasks by introducing the new task of Object State Change Anticipation in videos. OSCA focuses on the multi-class prediction of the state change occurring on an object during the next, yet unseen at inference time, action during a long procedural activity. Specifically, as shown in Fig. 1, at a certain decision point in time that is at the start of the next, yet unobserved action, we aim to predict the object state change class that will occur. The object state change will occur at the “Point of No Return” timestamp during the next action [19].
Figure 3: The intricate relation between verb/object/action and object state change. From left to right: one verb may signify different state changes; different verbs might signify the same state change; an action might lead to a variety of object state changes.
The OSCA task differs significantly from action anticipation, as it focuses on predicting imminent changes in the object’s state, if any, without requiring the prediction of the verb and possibly the noun categories associated with the anticipated state-modifying action. Although the OSCA task involves fewer target state change categories than action categories, it remains challenging. A single action may apply to various objects and contexts, resulting in a range of possible state changes, as shown in Fig. 2 and 3. In addition, while it might seem that, given a prediction for the next action, determining the next object state change would be straightforward, estimating such predictions in instructional videos is a challenging task, as current benchmarks suggest. Based on the Ego4D leaderboard for short-term action anticipation benchmark [^2] https://eval.ai/web/challenges/challenge-page/1623/leaderboard/3910/Overall (Top5 mAP: Overall approaches score 7%, verb and noun combined scores close to 17%, and noun-only score around 37% as of 11/2024), it is clear that a lot of research effort is still needed to devise efficient solutions to action anticipation in procedural activities and consequently their contribution to the estimation of the next state of the action-participating objects.
@@
To tackle this new task, we introduce a new formulation leveraging on the cumulative history of the textual description of recognized preceding actions and object states, up to the decision point where this prediction is performed. The integration of the former information with visual information concerning the recent past is a key idea to model this historical context effectively. Our approach may also predict that no object state change will occur in the forthcoming action which implies the anticipation of a no-state-modifying action. We evaluate the proposed approach toward the newly proposed OSCA problem to establish baseline results for automated change state anticipation in long instructional videos and to also investigate the potential impact and effectiveness of the extracted information in other related vision-based anticipation tasks. We assess the performance of our proposed method for the OSCA problem and show initial results based on a proposed extension of the popular Ego4D video dataset. We hence build on the large-scale and challenging Ego4D dataset [19] which provides egocentric videos by augmenting the available annotation data with labels for the object state changes based on the initial and the final object states for any state-modifying actions in a subset of videos related to the Hand & Object Interactions benchmark [^3] https://ego4d-data.org/docs/benchmarks/hands-and-objects/. This results in the Ego4D-OSCA, a variant of the Ego4D dataset that will become available to the community. Thus, the main contributions are:
@@ 2 Related Work
Object states capture dynamic aspects of object appearance and/or functionality and are subject to visually perceivable changes, as a result of state-modifying actions. They are also known as object fluents related to changeable object attributes [26, 2, 27]. Since there is no prior work on the problem of Object State Change Anticipation that we introduce, we review the literature on the closely related topics of action and next-active object anticipation, object state classification in images as well as the interplay between object state estimation and action recognition in videos.
@@ 2.1 Object State Classification/Recognition
Object State Change Estimation & Action Recognition in Videos: Object state changes have been considered a meaningful information source in video-based human action understanding and recognition (HAR). In HAR, object state changes are often considered complementary attributes to the visual representation of actions. These changes are typically derived within the visual domain through the utilization of explicit models for object detection and state estimation [11, 44, 28], or indirect modeling of object states based on general scene changes resulting from action execution [1, 5, 4]. Several methods exploit object states implicitly to estimate the type of action performed. The work in [1] was among the first to propose a method to automatically discover object states and the associated manipulation actions from videos by leveraging a discriminative clustering framework that jointly models the temporal order of object states and manipulation actions. The work in [27] explored the recognition of object fluents (changeable object attributes) and tasks (goal-oriented human activities) in egocentric videos using a hierarchical model that represents tasks as concurrent and sequential object fluents. Moreover, [29] focuses on understanding human actions within videos by analyzing complex interactions across multiple interrelated objects by recognizing their different state changes. In [46, 44] a multi-task self-supervised framework is proposed that allows the temporal localization of object state changes and state-modifying actions in uncurated web videos.
@@ 2.2 Action & Next-Active Object Anticipation
Figure 4: Two of the nine object state change super-annotated classes in the Ego4D-OSCA dataset, ‘deposit’ and ‘remove’. Pre-/post-state labels for these actions are shown as distinct video segments. ‘Deposit’ and ‘remove’ are inverse changes, where pre-deposit matches post-remove, and pre-remove matches post-deposit, indicated by frames and shapes of the same color.
@@ 3 Ego4D-OSCA Dataset
Dataset annotation: To enrich the LTA benchmark, we integrate object state annotations extracted from the dataset as follows. The original Ego4D dataset does not include annotations for the specific state labels of individual video frames. Instead, annotations about state changes are provided, which relate to entire video segments. Additionally, the dataset includes annotations for bounding boxes and object classes across seven critical frames within each video segment. These frames are temporally centered around the occurrence of the state change that occurs within each video segment. Based on this information, we super-annotate certain critical frames of each video segment with state-related labels as follows. For each video segment, we annotate the initial and final frames as $pre\_X$ and $post\_X$, respectively, where $X$ denotes the label of the state change. Furthermore, in line with the semantic implications of these changes, we establish three pairs of state changes. Each pair is constructed under the premise that the first action is the inverse of the second concerning the resulting state change. For instance, if $X$ and $Y$ represent inverse state changes, then the labels $pre\_X$ and $post\_Y$ are considered samples of identical states. A similar correspondence applies between $pre\_Y$ and $post\_X$. For example, the states $pre\_remove$ and $post\_deposit$ are considered identical, since $remove$ and $deposit$ constitute a pair of inverse state changes. Figure 4 delineates the specifics for these two super-annotated state change classes. The same condition is true for the state change classes of activate-deactivate and construct-deconstruct. The full set of pre- and post- object state pairs that constitute the target set of object state changes appear in the supplementary material. It is important to note that there exists a subset of video segments containing actions that do not induce state changes and, therefore, these segments are not considered for annotation.
@@
Details on the annotation process: The annotation process for the state transitions is applied to the $pre\_Y$ and $post\_X$ frames in each video segment. Overall, the annotation process consists of the next 4 steps (a schematic representation of the annotation pipeline is shown in Fig. 5). First, the PNR moment of the video segment being examined for annotation is compared to the PNR moment of the segment that has been previously annotated. If the PNR of the previously annotated segment is located after the PNR of the segment under examination, then the segment under examination is rejected. The reasoning behind this decision has to do with the learning of the segment features related to state transitions. This alignment of the two PNRs signifies that there is an overlap between the state transition actions of the two segments and therefore the feature learning becomes more challenging. Subsequently, it is examined if the object undergoing state change is occluded. If this is the case, the frame is rejected. Then, the bounding box area of the object is evaluated, and if it is below 100 square pixels—a threshold empirically chosen and commonly used in annotation tools like Voxel-51—the frame is discarded. Finally, the frame that has passed all the previous checks is annotated with the appropriate state label that pertains to the state transition action.
Dataset Statistics: The proposed Ego4D-OSCA dataset is compiled using a subset of the popular, large-scale Ego4D v2 dataset [19] that contains egocentric videos for a large variety of human daily living or work activities. In Table 1 we compare the proposed Ego4D-OSCA dataset with existing image and video datasets that also provide annotations related to object states. Ego4D-OSCA contains 61.858 training and 31.846 testing clips. The target tasks performed using each of the datasets are also noted. More dataset statistics can be found in the supplementary material.
@@ 4 Object State Change Anticipation - Baseline
The proposed framework, depicted in Fig. 6, draws inspiration from the efficacy of combining visual and lexical information for semantic action/activity encoding. To achieve this, it adopts a three-stream architecture. Within this design, a visual encoding module is tasked with capturing the visual attributes of ongoing actions, while two lexical-based encoders are employed to extract the semantic nuances from a procedural-oriented representation of past actions and object states. The framework fuses these distinct representations towards the unified objective of anticipating the next object state. This task entails the estimation of the forthcoming state in which the object of interest will reside during the subsequent action. The framework tries to holistically capture the underlying dynamics and contextual intricacies governing object-state transformations across sequential actions by integrating visual and lexical cues.
@@
Action & State History Encoders: As illustrated in Fig. 6, both encoders exploit the model design of the lexical encoder of the VLMAH model [31], which follows a NLP neural network design consisting of BiLSTM and MLP components. The decision to employ a simple NN for encoding the history, instead of utilizing LLMs was motivated by several factors. Firstly, the computational efficiency of LLMs such as GPT- or LLaMA, often entails significant resource requirements for training and inference [52], whereas a simpler neural network architecture mitigates computational overhead. Secondly, LLMs are pre-trained on general text corpora and may not capture the domain-specific nuances inherent in the textual data related to action histories and object states. Additionally, the simplicity of the chosen architecture facilitates interpretability, data efficiency, and customization, affording greater control over the model’s behavior and adaptation to the task’s requirements.
@@
During training, the proposed framework leverages oracle action and state detectors to provide the action and state history, respectively, for each clip in the dataset (see Fig. 6). These detectors estimate the current action and object state observed in the clip, serving as ground truth annotations for training purposes. However, it is important to note that for inference toward real-world applications, there is a requirement for current action and object state recognition models to provide input to the framework. Consequently, our model is solely tasked with the learning objective of next-object anticipation, focusing exclusively on predicting the future state of the object. By decoupling the training and inference phases in this manner, the model can effectively learn the dynamics of object-state transitions without the added complexity of simultaneously predicting the current action.
@@ 5 Implementation, Experiments and Results
Evaluation metrics: The evaluation of all examined models was conducted using top-1/5 mean accuracy, and F-score, following standard practices in the relevant literature.
@@ 5.1 OSCA Results
OSCA under actual action and state recognition: We conducted experiments where the oracles (action recognizer, object’s current state estimator), were substituted with existing baseline models. In this setting, the output of these models is utilized to populate the action and state histories that OSCA utilizes. We employed the following models: (a) the well-established SlowFast [12] model for action recognition; (b) the object-agnostic state classification method proposed by [16] as the current state classifier, with minor modifications, as described below.
@@
Action recognizer: Regarding the action classification module, we fine-tuned the SlowFast model [12] on a subset of the original Ego4D dataset. Specifically, we obtained the train/validation/test splits using the training set provided for the (LTA) long-term anticipation task of Ego4D based on the $60/20/20$ split scheme. The adaptation of the LTA data to the action recognition task resulted in $5754$ action classes and a total of $\approx 65K$ video clips (with a mean of $\approx 10.7$ sample clips per action class). SlowFast achieved $12.86\%$ Top-1 and $33.69\%$ Top-5 accuracy for the task of current action recognition. This low performance can be attributed to the extensive number of action classes and limited samples per class, as well as in Ego4D’s inherent motion and appearance similarities across different actions, e.g. take cup - take bottle, tie string - tie rope.
Model Top@1/5 mAcc F1-score
VID-A 23.93 / 89.10% 11.74%
VNLP (O-Action) 25.59 / 83.06% 24.62%
VNLP (O-State) 40.07 / 90.83% 33.57%
VNLP (O-Action, O-State) 39.20 / 89.76% 37.12%
VNLP (Action [12]) 23.04 / 81.31% 22.09%
VNLP (State [16]) 32.72 / 92.16% 21.78%
VNLP (Action [12], State [16]) 29.42 / 94.65% 26.29%
@@
Table 2: OSCA performance for various model configurations (O-: Oracle recognizer, VID-A: vision-only state anticipation model).
@@
Object state recognizer: Additionally, for the object-agnostic state history, we adapt the model of [16]. This model relies on the outputs of two distinct state classifiers. Each classifier receives the first (pre) or the last (post) frame of each video segment as input to predict the object state label for the respective frame. The prediction of the state-change label for the video segment considers both outputs and is derived based on the following rules. If the object state predictions are $pre\_X$ and $post\_X$, respectively, the inferred state change for the video segment is denoted as $X$. Conversely, if the classifiers predict $pre\_X$ and $post\_Y$, where $X$ and $Y$ are distinct and represent inverse state changes, it is concluded that no state change has occurred. Finally, if neither of the above conditions is met, the prediction of the state change defaults to the output of the second classifier; that is, if the prediction is $post\_Y$, the state change for the video segment is identified as Y. For example, if the predictions of the two classifiers are $pre\_activate$ and $post\_activate$ the prediction of state change would be $activate$. Likewise, if the predictions are $pre\_activate$ and $post\_deactivate$ the prediction of the state change would be that of no change. This object-agnostic state recognizer showcased $25.4\%$ mean state recognition accuracy.
In our experiments, replacing oracles with realistic recognizers to populate the action and state history buffers that are considered by the baseline OSCA model, we observed a significant accuracy drop (last block of lines in Table 2). This accuracy difference underscores the critical role of precise recognition of the current action and object state for effective anticipation of near-future object states within dynamic environments. Given the class imbalance in the proposed Ego4D-OSCA data set, the F1 score is a more appropriate performance measure. This metric considers both precision and recall while remaining insensitive to the true negatives of majority classes, unlike accuracy, which can be biased toward the majority class. Based on this rationale, the reported F1-scores in Table 2 indicate that a model combining both action and state history (past context) may be more effective for the object state anticipation task.